# Quality-Aware Image-Text Alignment for Real-World Image Quality Assessment

### v2

@@ Abstract
No-Reference Image Quality Assessment (NR-IQA) focuses on designing methods to measure image quality in alignment with human perception when a high-quality reference image is unavailable. The reliance on human-annotated Mean Opinion Score (MOS) in the majority of state-of-the-art NR-IQA approaches limits their scalability and broader applicability to real-world scenarios. To overcome this limitation, we propose QualiCLIP (Quality-aware CLIP), a CLIP-based self-supervised opinion-unaware method that does not require MOS. In particular, we introduce a quality-aware image-text alignment strategy to make CLIP generate quality-aware image representations. Starting from pristine images, we synthetically degrade them with increasing levels of intensity. Then, we train CLIP to rank these degraded images based on their similarity to quality-related antonym text prompts. At the same time, we force CLIP to generate consistent representations for images with similar content and the same level of degradation. Our method significantly outperforms other opinion-unaware approaches on several datasets with authentic distortions. Moreover, despite not requiring MOS, QualiCLIP achieves state-of-the-art performance even when compared with supervised methods in cross-dataset experiments, thus proving to be suitable for application in real-world scenarios. The code and the model are publicly available at https://github.com/miccunifi/QualiCLIP.
@@ 1 Introduction
Image Quality Assessment (IQA) aims to automatically evaluate the quality of images in accordance with human judgments, represented by a Mean Opinion Score (MOS). Specifically, No-Reference IQA (NR-IQA) focuses on developing methods that do not require a high-quality reference image and that are consequently more easily applicable in real-world scenarios. NR-IQA plays a critical role in diverse industries and research domains. For example, given the large number of photos that are captured and shared daily on social media platforms, it is imperative to design approaches that can measure image quality objectively to be able to store and process these images effectively. However, for such approaches to be reliable, they need to exhibit strong generalization capabilities.
Most NR-IQA methods are opinion-aware, i.e. they require human-labeled MOS as supervision during the training process [37, 31, 51, 52]. Some approaches, such as HyperIQA [37] or LIQE [51], directly train the model parameters on IQA datasets. Other methods, namely QPT [52] or Re-IQA [31], pre-train an encoder on unlabeled data via self-supervised learning and then either fine-tune the encoder weights or train a linear regressor using MOS. However, annotating IQA datasets is expensive and resource-intensive, as several human ratings are needed for each image for its MOS to be reliable. For example, the FLIVE dataset [47], which contains 40K real-world images, required about 4M ratings, up to 50 for a single image. The need for human annotations significantly hinders the scalability of opinion-aware approaches. In addition, these methods show limited generalization capabilities and thus applicability to real-world scenarios, as their performance considerably deteriorates on unseen datasets. To remove the requirement for expensive MOS, several opinion-unaware approaches have been proposed [23, 3, 10]. Nevertheless, existing opinion-unaware methods achieve significantly lower performance than supervised approaches in cross-dataset experiments, thus exhibiting limited applicability.
@@
To address this issue, we propose a quality-aware image-text alignment strategy that relies on self-supervised learning to remove the need for human annotations. We start by synthetically degrading pairs of pristine images using increasing levels of intensity. Then, we measure the similarity between each image and antonym prompts related to image quality, such as “Good photo” and “Bad photo”. We refer to these prompts as positive and negative prompts, respectively. Finally, we employ a training strategy based on a margin ranking loss [15, 18] that allows us to achieve two objectives. First, we want CLIP to generate consistent representations for images having similar content and comparable quality, i.e. exhibiting the same amount of distortion. Second, the similarity between the positive (negative) prompt and the increasingly degraded versions of the images must correlate inversely (directly) with the intensity of the distortion. Our approach, named QualiCLIP (Quality-aware CLIP), is both self-supervised and opinion-unaware, as we do not rely on any form of supervision – especially MOS – at any step of the training process. Thanks to our training strategy, the image-text alignment in CLIP’s embedding space focuses more on low-level image characteristics than high-level semantics. Consequently, QualiCLIP generates image representations whose similarity to the antonym prompts correlates with the inherent quality of the images, as illustrated in Fig. 1.
The experiments show that the proposed approach obtains significant performance improvements – up to a 20% relative gain – over existing opinion-unaware methods on several datasets with authentic distortions. Furthermore, QualiCLIP achieves state-of-the-art results even in comparison with supervised techniques in the cross-dataset setting, thus exhibiting suitability for real-world applications. Finally, the gMAD [22] competition and visualization with gradCAM [32] show that QualiCLIP demonstrates both greater robustness and enhanced explainability than competing methods.
@@
  • Our method significantly outperforms existing opinion-unaware approaches and achieves state-of-the-art results even when compared to supervised techniques in cross-dataset experiments.
@@ 2 Related Work
No-reference image quality assessment Due to its wide range of applications in real-world scenarios, in recent years research on NR-IQA has gained significant momentum [26, 37, 24, 1, 3]. Traditional methods [26, 25, 49] rely on extracting hand-crafted image features to derive quality scores. Such approaches achieve promising performance on synthetic datasets but struggle on images with authentic distortions. More recently, several methods relying on supervised learning have been introduced [37, 8, 24, 1, 46, 31, 34]. Some approaches directly employ MOS during model training [37, 8, 46]. For example, HyperIQA [37] proposes a self-adaptive hypernetwork that separates content understanding from quality prediction. Another research direction involves pre-training an encoder on unlabeled images via self-supervised learning. Then, the image representations are mapped to quality scores by fine-tuning the encoder weights [52] or training a linear regressor [24, 31, 1] on human annotations. For instance, QPT [52] and Re-IQA [31] use a contrastive loss to train the encoder to separate between images degraded with different types and degrees of distortion. Despite their impressive performance, the scalability and applicability of supervised methods are limited by their need for costly human annotations. This requirement is removed by opinion-unaware approaches [26, 49, 23, 3, 35]. Some of them, such as NIQE [26], are based on natural scene statistics [26, 49], while others employ self-supervised learning [23, 3, 35, 10]. For example, CL-MI [3] pre-trains an encoder on synthetic data and then fine-tunes it on authentic images via a mutual information-based loss. Nevertheless, existing opinion-unaware approaches fall behind supervised methods in cross-dataset experiments. In contrast, despite not requiring MOS, our method achieves state-of-the-art performance on unseen datasets even when considering supervised techniques.
@@ 3.2 Synthetic Degradation with Increasing Levels of Intensity
Several works have shown the effectiveness of training self-supervised NR-IQA models on synthetically degraded images [24, 52, 31, 1]. Since the synthetic distortions are designed to resemble real-world ones, such models manage to achieve competitive performance also on authentic datasets. Therefore, we propose to synthetically degrade unlabeled pristine images using progressively higher levels of intensity. In this way, we can train our model in a self-supervised way to rank the different versions of each image according to the severity of their degradation. Following [1], we consider 24 distinct degradation types spanning the 7 distortion groups defined by the KADID-10k dataset [17]: 1) Brightness change; 2) Blur; 3) Spatial distortions; 4) Noise; 5) Color distortions; 6) Compression; 7) Sharpness & contrast. Each distortion has $L\!=\!5$ levels of intensity. See the supplementary material for more details on the specific degradation types. Fig. 2 shows some examples of degraded images for varying degrees of intensity.
@@ 3.3 Quality-Aware Image-Text Alignment
where $\tau$ is a temperature hyperparameter. Note that, since we keep CLIP’s text encoder weights frozen, we need to compute the text features of the antonym prompts only once, and we can use them both for training and inference. Therefore, at inference time the computational cost of our method is the same as an image-encoder-only model with the same backbone.
BID KonIQ-10k CLIVE FLIVE SPAQ Average
Method SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
NIQE [26] 0.458 0.460 0.526 0.534 0.450 0.493 0.158 0.221 0.703 0.712 0.459 0.484
IL-NIQE [49] 0.495 0.512 0.493 0.519 0.438 0.503 0.165 0.209 0.710 0.717 0.460 0.492
CONTRIQUE-OU [24] 0.564 0.552 0.637 0.630 0.394 0.422 0.199 0.228 0.676 0.680 0.494 0.502
Re-IQA-OU [31] 0.395 0.370 0.558 0.550 0.418 0.444 0.218 0.238 0.616 0.618 0.441 0.444
ARNIQA-OU [1] 0.536 0.515 0.741 0.760 0.484 0.558 0.299 0.362 0.789 0.797 0.570 0.598
CL-MI [3] 0.428 0.395 0.645 0.645 0.507 0.525 0.257 0.293 0.701 0.702 0.508 0.512
CLIP-IQA [41] 0.634 0.645 0.699 0.733 0.611 0.593 0.287 0.349 0.733 0.728 0.593 0.610
GRepQ-OU [36] 0.575 0.553 0.768 0.788 0.740 0.769 0.327 0.438 0.805 0.809 0.643 0.671
QualiCLIP 0.646 0.657 0.817 0.841 0.756 0.799 0.392 0.499 0.848 0.857 0.692 0.731
Table 1: Comparison between QualiCLIP and competing opinion-unaware methods. Best and second-best scores are highlighted in bold and underlined, respectively. OU indicates Opinion-Unaware version, as detailed in Sec. 4.1.
@@
Discussion Our quality-aware image-text alignment strategy stems from the inherent limitations of applying common self-supervised NR-IQA training techniques to CLIP. Prior methods, such as QPT [52] and Re-IQA [31], train an encoder using a contrastive loss that maximizes the similarity between the representations of crops from the same degraded image, while minimizing the similarity with the representations of crops coming from different images within the same batch. While this strategy proved its effectiveness for image-encoder-only models, applying it to CLIP’s image encoder would introduce a significant mismatch with CLIP’s training process [28]. Indeed, CLIP is trained using an inter-modal (i.e. image-text) objective, aligning the features – extracted with the image and text encoders – of corresponding images and texts within a common embedding space. Consequently, fine-tuning CLIP’s image encoder by considering only intra-modal (i.e. image-image) similarities without exploiting its alignment with the text encoder contradicts its design. For this reason, we propose to train our model by employing image-text similarities to leverage CLIP’s inherent inter-modal alignment. Additionally, using a contrastive loss to maximize (or minimize) the similarity of the antonym prompts with multiple different training samples within the same batch would correspond to making assumptions on the relative quality of unlabeled images with completely different content, which is unfeasible [52]. Instead, by relying on a ranking loss that only considers progressively degraded versions of the same image, we can leverage their inherent quality ranking as supervision to train our model in an effective way.
@@ 4.1 Quantitative Results
Evaluation protocol We evaluate the performance using Spearman’s Rank-order Correlation Coefficient (SRCC) and Pearson’s Linear Correlation Coefficient (PLCC), which measure prediction monotonicity and accuracy, respectively. Higher values of SRCC and PLCC correspond to better results. Following [9], we pass the quality predictions through a four-parameter logistic non-linearity before computing PLCC. We test our method on five authentic datasets, each comprising real-world images annotated with human judgments of picture quality in the form of a MOS. Specifically, we consider: BID [5], KonIQ-10k [12], CLIVE [7], FLIVE [47], and SPAQ [6]. See the supplementary material for more details on the datasets. We employ the whole dataset for testing.
@@
We compare our approach to state-of-the-art methods in two settings: zero-shot and cross-dataset. Our method remains consistent across both settings; the only variation lies in the competing methods. For a fair comparison, we compute the results of the baselines using our evaluation protocol. For each baseline, we employ the official pre-trained model when available, or, otherwise, train the model following the procedure described in the original paper. In the zero-shot setting, we only consider opinion-unaware methods [26, 49, 3, 41] and approaches that, with slight modifications, can function without requiring MOS [24, 31, 1, 36]. In particular, we follow the original paper for GRepQ [36], while for methods based on a linear regressor [24, 31, 1], such as CONTRIQUE [24], we employ a NIQE-style framework on the extracted image features, similar to [3]. In the cross-dataset setting, we compare with supervised methods [37, 8, 24, 31, 1, 41, 36] using testing datasets different from the training one, simulating real-world scenarios. This experiment allows us to compare the generalization capabilities of our model with supervised approaches. We employ FLIVE as the training dataset for the baselines. For a fair comparison, we train LIQE [51] using a ResNet50 backbone and restrict our analysis to methods that do not rely on an MLLM, as these models entail substantial computational demands.
Opinion BID KonIQ-10k CLIVE SPAQ Average
Method unaware SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
HyperIQA [37] 0.590 0.579 0.738 0.742 0.736 0.743 0.653 0.658 0.679 0.681
TReS [8] 0.637 0.638 0.748 0.751 0.735 0.751 0.743 0.741 0.716 0.720
CONTRIQUE [24] 0.585 0.599 0.779 0.781 0.734 0.751 0.817 0.825 0.729 0.739
Re-IQA [31] 0.627 0.635 0.789 0.825 0.719 0.770 0.826 0.830 0.740 0.765
ARNIQA [1] 0.635 0.638 0.787 0.804 0.734 0.777 0.841 0.849 0.749 0.767
CLIP-IQA+ [41] 0.642 0.654 0.784 0.801 0.707 0.732 0.751 0.750 0.721 0.734
LIQE [51] 0.613 0.603 0.806 0.807 0.749 0.770 0.813 0.816 0.745 0.749
GRepQ [36] 0.630 0.608 0.807 0.812 0.768 0.785 0.828 0.836 0.758 0.760
QualiCLIP 0.646 0.657 0.817 0.841 0.756 0.799 0.848 0.857 0.767 0.789
Table 2: Comparison between QualiCLIP and supervised methods trained on the FLIVE [47] dataset. Best and second-best scores are highlighted in bold and underlined, respectively.
Zero-shot setting We report the results for the zero-shot setting in Tab. 1. Our approach achieves state-of-the-art performance on all the testing datasets, consistently outperforming the other methods by a significant margin. In particular, QualiCLIP obtains a relative gain of about 20% and 7% over the best-performing baseline on the FLIVE dataset and on average, respectively. The improvement over CLIP-IQA proves that the proposed method generates more accurate quality-aware image representations compared to off-the-shelf CLIP models. In addition, we recall that GRepQ employs a low-level encoder and a high-level fine-tuned CLIP-based encoder. Despite relying solely on CLIP, QualiCLIP outperforms GRepQ, further confirming the effectiveness of our approach.
Cross-dataset setting Tab. 2 shows the results for the cross-dataset setting. Despite not requiring MOS, QualiCLIP achieves state-of-the-art performance across all testing datasets, surpassing the baselines in most scenarios. This consistently strong performance suggests that our model is well-suited for real-world applications, where obtaining robust and reliable results is crucial. Finally, comparing Tabs. 1 and 2, we observe that QualiCLIP stands out as the only opinion-unaware approach to achieve state-of-the-art performance even against supervised methods.

### v3

# Quality-Aware Image-Text Alignment for Opinion-Unaware Image Quality Assessment

@@ Abstract
No-Reference Image Quality Assessment (NR-IQA) focuses on designing methods to measure image quality in alignment with human perception when a high-quality reference image is unavailable. Most state-of-the-art NR-IQA approaches are opinion-aware, i.e. they require human annotations for training. This dependency limits their scalability and broad applicability. To overcome this limitation, we propose QualiCLIP (Quality-aware CLIP), a CLIP-based self-supervised opinion-unaware approach that does not require human opinions. In particular, we introduce a quality-aware image-text alignment strategy to make CLIP generate quality-aware image representations. Starting from pristine images, we synthetically degrade them with increasing levels of intensity. Then, we train CLIP to rank these degraded images based on their similarity to quality-related antonym text prompts. At the same time, we force CLIP to generate consistent representations for images with similar content and the same level of degradation. Our experiments show that the proposed method improves over existing opinion-unaware approaches across multiple datasets with diverse distortion types. Moreover, despite not requiring human annotations, QualiCLIP achieves excellent performance against supervised opinion-aware methods in cross-dataset experiments, thus demonstrating remarkable generalization capabilities. The code and the model are publicly available at https://github.com/miccunifi/QualiCLIP.
@@ 1 Introduction
Image Quality Assessment (IQA) aims to automatically evaluate the quality of images in accordance with human judgments represented by a Mean Opinion Score (MOS). Specifically, No-Reference IQA (NR-IQA) focuses on developing methods that do not require a high-quality reference image and that are consequently more easily applicable in real-world scenarios. NR-IQA plays a critical role in diverse industries and research domains. For example, given the large number of photos that are captured and shared daily on social media platforms, it is imperative to design approaches that can measure image quality objectively to store and process these images effectively.
Most NR-IQA methods are opinion-aware, i.e. they require human-labeled MOS as supervision during the training process [42, 35, 58, 60]. Some approaches, such as HyperIQA [42] or LIQE [58], directly train the model parameters on IQA datasets. Other methods, namely QPT [60] or Re-IQA [35], pre-train an encoder on unlabeled data via self-supervised learning and then either fine-tune the encoder weights or train a linear regressor using MOS. However, annotating IQA datasets is expensive and resource-intensive, as several human ratings are needed for each image for its MOS to be reliable. For example, the FLIVE dataset [53], which contains 40K real-world images, required about 4M ratings, up to 50 for a single image. The need for human annotations significantly hinders the scalability of opinion-aware approaches. In addition, these methods show limited generalization capabilities and thus applicability to scenarios where labeled data is unavailable, as their performance considerably deteriorates on unseen datasets. To remove the requirement for expensive MOS, several opinion-unaware approaches have been proposed [25, 3, 9]. For instance, CL-MI [3] introduces a two-stage self-supervised approach that employs two different training strategies for synthetically and authentically degraded images. Nevertheless, existing opinion-unaware methods achieve significantly lower performance than opinion-aware approaches in cross-dataset experiments, thus exhibiting limited applicability.
@@
To address this issue, we propose a quality-aware image-text alignment strategy that relies on self-supervised learning to remove the need for human annotations. We start by synthetically degrading pairs of pristine images using increasing levels of intensity. Then, we measure the similarity between each image and antonym prompts related to image quality, such as “Good photo” and “Bad photo”. We refer to these prompts as positive and negative prompts, respectively. Finally, we employ a training strategy based on a margin ranking loss [16, 19] that allows us to achieve two objectives. First, we want CLIP to generate consistent representations for images having similar content and comparable quality, i.e. exhibiting the same amount of distortion. Second, the similarity between the positive (negative) prompt and the increasingly degraded versions of the images must correlate inversely (directly) with the intensity of the distortion. Our approach, named QualiCLIP (Quality-aware CLIP), is both self-supervised and opinion-unaware, as we do not rely on any form of supervision – especially MOS – at any step of the training process. Thanks to our training strategy, the image-text alignment in CLIP’s embedding space prioritizes low-level image characteristics over high-level semantics. Consequently, QualiCLIP generates image representations whose similarity to the antonym prompts correlates with the inherent quality of the images, as illustrated in Fig. 1.
The experiments demonstrate that the proposed approach improves over existing opinion-unaware methods across multiple datasets encompassing various degradations. Furthermore, QualiCLIP is the only opinion-unaware approach that consistently obtains remarkable results even when compared against supervised opinion-aware techniques in the cross-dataset setting. The strong and robust performance of our model across different datasets highlights its commendable generalization capabilities and suitability for real-world applications.
@@
  • Our method improves on existing opinion-unaware approaches across multiple datasets and achieves excellent results even when compared to supervised opinion-aware techniques in cross-dataset experiments.
@@ 2 Related Work
No-Reference Image Quality Assessment Due to its wide range of applications in real-world scenarios, in recent years research on NR-IQA has gained significant momentum [29, 42, 26, 1, 3]. Traditional methods [29, 28, 55] rely on extracting hand-crafted image features to derive quality scores. Such approaches achieve promising performance on synthetic datasets but struggle on images with authentic distortions. More recently, several methods relying on supervised learning have been introduced [42, 7, 26, 1, 52, 35, 38]. Some approaches directly employ MOS during model training [42, 7, 52]. For example, HyperIQA [42] proposes a self-adaptive hypernetwork that separates content understanding from quality prediction. Another research direction involves pre-training an encoder on unlabeled images via self-supervised learning. Then, the image representations are mapped to quality scores by fine-tuning the encoder weights [60] or training a linear regressor [26, 35, 1] on human annotations. For instance, QPT [60] and Re-IQA [35] use a contrastive loss to train the encoder to separate between images degraded with different types and degrees of distortion. Despite their impressive performance, the scalability and applicability of supervised methods are limited by their need for costly human annotations. This requirement is removed by opinion-unaware approaches [29, 55, 25, 3, 40, 30]. Some of them, such as NIQE [29], are based on natural scene statistics [29, 55], while others employ self-supervised learning [25, 3, 40, 9]. For example, CL-MI [3] pre-trains an encoder on synthetic data and then fine-tunes it on authentic images via a mutual information-based loss. Nevertheless, existing opinion-unaware approaches fall behind supervised methods in cross-dataset experiments. In contrast, despite not requiring MOS, our method achieves remarkable performance on unseen datasets even when considering opinion-aware techniques.
@@ 3.2 Synthetic Degradation with Increasing Levels of Intensity
Although authentic distortions cannot be perfectly replicated synthetically, prior studies have shown that synthetic distortions remain effective for training self-supervised NR-IQA models that generalize well to real-world images [26, 60, 35, 1]. Following these works, we synthetically degrade unlabeled pristine images using progressively increasing levels of intensity. In this way, we can train our model in a self-supervised way to rank the different versions of each image according to the severity of their degradation. Following [1], we consider 24 distinct degradation types spanning the 7 distortion groups defined by the KADID-10k dataset [18]: 1) Brightness change; 2) Blur; 3) Spatial distortions; 4) Noise; 5) Color distortions; 6) Compression; 7) Sharpness & contrast. Each distortion has $L\!=\!5$ levels of intensity. See the supplementary material for more details on the specific degradation types. Fig. 2 shows some examples of degraded images for varying degrees of intensity.
@@ 3.3 Quality-Aware Image-Text Alignment
where $\tau$ is a temperature hyperparameter. Note that, since we keep CLIP’s text encoder weights frozen, we need to compute the text features of the antonym prompts only once, and we can use them both for training and inference. Therefore, at inference time, the computational cost of our method is the same as that of an image-encoder-only model with the same backbone.
Authentic Datasets Image Restoration Datasets AIGC Datasets
KonIQ-10k CLIVE FLIVE SPAQ CVIU PIPAL AGIQA-1K AGIQA-3K
Method OU SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
CONTRIQUE [26] 0.874 0.882 0.806 0.838 0.596 0.648 0.910 0.912 0.912 0.916 0.581 0.611 0.799 0.855 0.817 0.879
Re-IQA [35] 0.883 0.887 0.783 0.797 0.623 0.685 0.909 0.912 0.887 0.895 0.568 0.587 0.783 0.84 0.811 0.874
ARNIQA [1] 0.869 0.883 0.797 0.823 0.595 0.670 0.904 0.909 0.919 0.926 0.634 0.666 0.768 0.849 0.803 0.881
CLIP-IQA+ [47] 0.873 0.890 0.815 0.849 0.602 0.662 0.901 0.904 0.925 0.918 0.552 0.558 0.817 0.855 0.844 0.894
GRepQ [41] 0.882 0.883 0.793 0.813 0.576 0.611 0.902 0.903 0.882 0.857 0.554 0.568 0.740 0.797 0.807 0.858
NIQE [29] 0.527 0.532 0.446 0.469 0.203 0.262 0.700 0.712 0.685 0.696 0.167 0.181 0.623 0.721 0.510 0.526
IL-NIQE [55] 0.511 0.532 0.427 0.481 0.234 0.286 0.713 0.724 0.595 0.668 0.231 0.220 0.645 0.757 0.528 0.544
CL-MI [3] 0.664 0.653 0.494 0.480 0.290 0.322 0.701 0.701 0.433 0.481 0.281 0.282 0.474 0.621 0.591 0.665
MDFS [30] 0.739 0.746 0.427 0.467 0.360 0.407 0.805 0.812 0.702 0.742 0.255 0.262 0.734 0.818 0.697 0.723
CONTRIQUE-OU [26] 0.651 0.637 0.348 0.364 0.321 0.341 0.677 0.685 0.610 0.715 0.225 0.218 0.644 0.717 0.465 0.489
Re-IQA-OU [35] 0.580 0.568 0.379 0.388 0.350 0.341 0.613 0.616 0.648 0.662 0.152 0.151 0.644 0.764 0.366 0.430
ARNIQA-OU [1] 0.746 0.762 0.468 0.557 0.428 0.468 0.788 0.797 0.673 0.712 0.327 0.331 0.667 0.786 0.614 0.679
CLIP-IQA [47] 0.695 0.730 0.656 0.670 0.344 0.473 0.733 0.734 0.613 0.642 0.274 0.279 0.498 0.651 0.638 0.711
GRepQ-OU [41] 0.776 0.798 0.722 0.766 0.289 0.423 0.802 0.805 0.748 0.785 0.434 0.419 0.161 0.530 0.613 0.734
QualiCLIP 0.817 0.838 0.725 0.802 0.442 0.556 0.841 0.851 0.840 0.852 0.410 0.424 0.736 0.805 0.667 0.735
[…]
@@
Discussion Our quality-aware image-text alignment strategy stems from the inherent limitations of applying common self-supervised NR-IQA training techniques to CLIP. Prior methods, such as QPT [60] and Re-IQA [35], train an encoder using a contrastive loss that maximizes the similarity between the representations of crops from the same degraded image, while minimizing the similarity with the representations of crops coming from different images within the same batch. While this strategy proved its effectiveness for image-encoder-only models, applying it to CLIP’s image encoder would introduce a significant mismatch with CLIP’s training process [32]. Indeed, CLIP is trained using an inter-modal (i.e. image-text) objective, aligning the features – extracted with the image and text encoders – of corresponding images and texts within a common embedding space. Consequently, fine-tuning CLIP’s image encoder by considering only intra-modal (i.e. image-image) similarities without exploiting its alignment with the text encoder contradicts its design [45, 27]. For this reason, we propose to train our model by employing image-text similarities to leverage CLIP’s inherent inter-modal alignment. Additionally, using a contrastive loss to maximize (or minimize) the similarity of the antonym prompts with multiple different training samples within the same batch would correspond to making assumptions on the relative quality of unlabeled images with completely different content, which is unfeasible [60]. Instead, by relying on a ranking loss that only considers progressively degraded versions of the same image, we can leverage their inherent quality ranking as supervision to train our model in an effective way.
Authentic Datasets Image Restoration Datasets AIGC Datasets
KonIQ-10k CLIVE SPAQ CVIU PIPAL AGIQA-1K AGIQA-3K
Method OU SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC SRCC PLCC
HyperIQA [42] 0.738 0.735 0.699 0.722 0.649 0.650 0.361 0.400 0.289 0.95 0.403 0.582 0.490 0.570
DBCNN [57] 0.804 0.823 0.660 0.698 0.808 0.817 0.786 0.803 0.356 0.374 0.525 0.726 0.510 0.657
TReS [7] 0.759 0.758 0.696 0.727 0.739 0.735 0.466 0.510 0.326 0.328 0.392 0.585 0.515 0.569
LIQE [58] 0.813 0.801 0.743 0.756 0.713 0.712 0.560 0.625 0.349 0.393 0.450 0.663 0.563 0.633
QCN [39] 0.728 0.789 0.653 0.731 0.815 0.819 0.753 0.757 0.370 0.382 0.458 0.705 0.518 0.595
CONTRIQUE [26] 0.777 0.777 0.734 0.766 0.820 0.827 0.616 0.681 0.357 0.390 0.485 0.663 0.464 0.540
Re-IQA [35] 0.796 0.827 0.690 0.761 0.825 0.830 0.714 0.741 0.306 0.302 0.406 0.675 0.554 0.625
ARNIQA [1] 0.798 0.810 0.699 0.765 0.837 0.846 0.632 0.673 0.351 0.361 0.437 0.682 0.487 0.585
CLIP-IQA+ [47] 0.783 0.798 0.694 0.741 0.755 0.754 0.747 0.770 0.319 0.332 0.502 0.595 0.505 0.518
GRepQ [41] 0.810 0.818 0.751 0.772 0.829 0.838 0.799 0.816 0.448 0.460 0.321 0.610 0.581 0.697
QualiCLIP 0.817 0.838 0.725 0.802 0.841 0.851 0.840 0.852 0.410 0.424 0.736 0.805 0.667 0.735
Table 2: Quantitative results of the cross-dataset evaluation setting. We employ the FLIVE [53] dataset to train the supervised methods. OU stands for Opinion-Unaware. Best and second-best scores are highlighted in bold and underlined, respectively.
@@ 4.1 Quantitative Results
We conduct several experiments to compare the performance of QualiCLIP with existing opinion-unaware and opinion-aware methods. In the supplementary material, we also study the robustness and explainability of our model.
@@
Evaluation protocol We evaluate the performance using Spearman’s Rank-order Correlation Coefficient (SRCC) and Pearson’s Linear Correlation Coefficient (PLCC), which measure prediction monotonicity and accuracy, respectively. Higher values of SRCC and PLCC correspond to better results. Following [8], we pass the quality predictions through a four-parameter logistic non-linearity before computing PLCC.
We evaluate our method on several IQA datasets, each containing images annotated with human judgments of picture quality in the form of MOS. These datasets feature images with various types of distortions, including authentic degradations, artifacts from image restoration methods, and AI-generated content. Specifically, we consider four authentic datasets: KonIQ-10k [11], CLIVE [6], FLIVE [53], and SPAQ [5]; two image restoration datasets: CVIU [22] and PIPAL [12]; and two AIGC datasets: AGIQA-1K [59] and AGIQA-3K [15]. Additional details on the datasets are provided in the supplementary material, where we also report experiments on images with synthetic distortions. Following [26, 1, 35], we randomly split the datasets into 70% for training, 10% for validation, and 20% for testing. For datasets that include reference images, namely the image restoration and synthetic datasets, we ensure that splits are made based on reference images to prevent content overlap. To mitigate selection bias in the training set, we repeat the training/testing process 10 times and report the median results. Due to its large size, for FLIVE, we follow [26, 1, 35] and use only the official train-validation-test split [53].
@@
We compare our approach to state-of-the-art methods in two settings: zero-shot and cross-dataset. Our method remains consistent across both settings; the only variation lies in the competing methods. For a fair comparison, we compute the results of the baselines using our evaluation protocol. For each baseline, we employ the official pre-trained model when available, or, otherwise, train the model following the procedure described in the original paper. In the zero-shot setting, we compare with existing opinion-unaware methods. In addition, following [3], we consider opinion-aware approaches that can be modified to function without requiring MOS (indicated with the suffix “-OU”). In particular, for GRepQ [41], we follow the zero-shot strategy detailed in the original paper. For methods based on an image encoder and a linear regressor, such as CONTRIQUE [26], we extract the image features via the pre-trained encoder and then employ a NIQE-style framework to predict the quality scores, similar to [3]. In the cross-dataset setting, we evaluate the generalization capabilities of our model by comparing it with supervised opinion-aware methods on testing datasets different from the training one. Due to its large scale, we employ FLIVE as the training dataset for the baselines. Additionally, we report results using CLIVE and PIPAL in the supplementary material. For a fair comparison, we train LIQE [58] using a ResNet50 backbone and restrict our analysis to methods that do not rely on an MLLM, as these models entail substantial computational demands.
Zero-shot setting We report the results for the zero-shot setting in Tab. 1. Our approach achieves the best performance on 13 out of 16 evaluation metrics and ranks second on the remaining 3, with SRCC improvements over the best-performing baseline of up to 9.2%, observed on the CVIU dataset. Notably, QualiCLIP sets the new state of the art for opinion-unaware approaches on authentic and image restoration datasets, proving the effectiveness of our training strategy. The improvement over CLIP-IQA highlights that our model generates more accurate quality-aware image representations than off-the-shelf CLIP models. Compared to GRepQ-OU, which is the strongest existing approach in most scenarios, the proposed method obtains better results on all the testing datasets excluding PIPAL. Moreover, while GRepQ-OU combines a low-level encoder with a high-level fine-tuned CLIP-based encoder, QualiCLIP relies solely on CLIP, making it more straightforward and efficient. For reference, Tab. 1 also includes the performance of supervised opinion-aware methods trained on the training split of each testing dataset. We observe that all opinion-unaware methods fall behind supervised opinion-aware approaches when a training set is available, showing that there is still room for improving the performance of opinion-unaware models.
Cross-dataset setting Tab. 2 shows the results for the cross-dataset setting. Despite not requiring MOS, QualiCLIP outperforms the baselines on 11 out of 14 evaluation metrics. Specifically, our method achieves excellent performance across datasets with various types of distortions, demonstrating its robustness. This makes our model well-suited for real-world applications where a training set is unavailable. Moreover, comparing Tabs. 1 and 2 reveals two key observations. First, the performance of supervised opinion-aware models significantly decreases when tested on unseen datasets (e.g. GRepQ on AIGC datasets), highlighting their limited generalization capabilities. Second, QualiCLIP stands out as the only opinion-unaware approach to consistently obtain remarkable results even against supervised opinion-aware methods.